OpenStudy (anonymous):

Write the point-slope equation of the line through (-2, 1) with a slope of 3. The answer I came up with was y - 2 = 3(x + 1)

OpenStudy (asnaseer):

not quite... remember in (-2, 1) the x-coordinate is -2 and the y coordinate is 1 also remember the point-slope form for a line passing through (x1, y1) is: y - y1 = m(x - x1) where m is the slope (as we saw in your previous question)

OpenStudy (asnaseer):

pay careful attention to the sign of the coordinates

OpenStudy (anonymous):

y - 1 = 3(x + 2)

OpenStudy (asnaseer):

yes - now next you need to simplify by expanding the right-hand-side and then taking the "-1" over to the right-hand-side as well